OpenStudy (xochital):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

first thing i always do with this type of problem is to make a chart

OpenStudy (anonymous):

tony = x Belinda = 8+2x making sense so far ?

OpenStudy (xochital):

yes so far it is

OpenStudy (anonymous):

ok so now we know that combined age is 56 so what do you think we do with the two to set up an equation?

OpenStudy (xochital):

t + b = 56 8 + 2t = b

OpenStudy (anonymous):

to set up an equation x+8+2x=56

OpenStudy (anonymous):

does that make sense ? because we have a total now combine like terms

OpenStudy (xochital):

oh okay so it would be 8 + 3x = 56 ??

OpenStudy (anonymous):

yes now we want to get the x onto one side

OpenStudy (xochital):

x = 16 right?

OpenStudy (anonymous):

do u have a solution or what the answer is supposed to be ?

OpenStudy (anonymous):

ok i goofed up so tony should be = 2x belinda = 8+2x

OpenStudy (xochital):

so tony's age is 16... so belinda must be 40

OpenStudy (xochital):

ohh okay

OpenStudy (anonymous):

ok sorry i did the addition wrong = you are right my bad :/

OpenStudy (xochital):

so tony is 16 and belinda is 40?

OpenStudy (anonymous):

yes !!

OpenStudy (xochital):

okay thank you :)